OpenStudy (anonymous):

Which of the following is the equation of a line parallel to the line y = 4x + 1, passing through the point (5,1)? A. 4x - y = 19 B. 4x + y = -19 C. -4x - y = 19 D. 4x + y = 19

OpenStudy (anonymous):

\[y-1=4(x-5)\]\[y-1=4x-20\]\[-4x+4y=-19\]\[4x-y=19\]answer A